### v1

@@ 5 Conclusion
We have constructed a class of PINN that is equipped with an uncertainty quantification framework for its target and PDE parameter inferences. It leverages the marginal distribution loss function of EDL for estimating uncertainty of outputs and infers unknown parameters of the PDE through learning a posterior distribution. Two key features distinguish our model from the usual PINN: (i)the loss weight of the PDE residual loss term evolves dynamically throughout model training, (ii)a prior distribution based on a purely data-fitted model regularizes the learning of the unknown PDE parameters. Validating our model on two controlled case studies involving the 1D Poisson equation with a Gaussian source and the 2D Fisher-KPP equation, we found that E-PINN generated empirical coverage probabilities that were calibrated significantly better than the B-PINN and Deep Ensemble methods. It inferred the unknown parameters accurately within the bounds of confidence intervals consistent with the Monte-Carlo goodness-of-fit test. In the Fisher-KPP case, E-PINN’s uncertainty distribution correlated well with a temporally varying noise, unlike the other two frameworks. As an illustrative real-world application, we demonstrated how E-PINN can be applied to model clinical IVGTT datasets presumed to follow Bergman minimal model of glucose-insulin dynamics, with results indicating that the insulin sensitivity index $S_{I}$ discriminates effectively between healthy and diabetic patients. We look forward to more extensive validation and application of E-PINN to solving inverse problems across scientific disciplines.

### v2

@@ 5 Conclusion
We have constructed a class of PINN that is equipped with an uncertainty quantification framework for its target and PDE parameter inferences. It leverages the marginal distribution-based loss function of EDL for estimating uncertainty of outputs and infers unknown parameters of the PDE through learning a posterior distribution. Two key features distinguish our model from the usual PINN: (i)the loss weight of the PDE residual loss term evolves dynamically throughout model training, (ii)a prior distribution based on a purely data-fitted model regularizes the learning of the unknown PDE parameters. Validating our model on two controlled case studies involving the 1D Poisson equation with a Gaussian source and the 2D Fisher-KPP equation, we found that E-PINN generated empirical coverage probabilities that were calibrated significantly better than the B-PINN and Deep Ensemble methods. It inferred the unknown parameters accurately within the bounds of confidence intervals consistent with the Monte-Carlo goodness-of-fit test. In the Fisher-KPP case, E-PINN’s uncertainty distribution correlated well with a temporally varying noise, unlike the other two frameworks. As an illustrative real-world application, we demonstrated how E-PINN can be applied to model clinical IVGTT datasets presumed to follow Bergman minimal model of glucose-insulin dynamics, with results indicating that the insulin sensitivity index $S_{I}$ discriminates effectively between healthy and diabetic patients. We look forward to more extensive validation and application of E-PINN to solving inverse problems across scientific disciplines.